OpenStudy (anonymous):

Solve the following equation. Be sure to check the answer in the original equation if you multiply both sides of the expression that contains the variable. (Problem in comments...)

OpenStudy (anonymous):

\[\frac{ 2x }{ x+8 } = \frac{ x }{ x+9 } -\frac{ 9 }{ x^2+17x+72 }\]

OpenStudy (anonymous):

2x/x+8= x/x+9 - 9/(x+8)(x+9)

OpenStudy (anonymous):

multiply fisrt variable with (x+9) and the second variable with (x+8)

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

\[\frac{2x ^{2}+18x }{ x^{2}+17x+72 } = \frac{ x^{2}+8x }{ x^{2}+17x+72 }-\frac{ 9 }{ x^{2}+17x+72 }\]

OpenStudy (anonymous):

\[\frac{ 2x^{2}+18x }{ x^{2}+17x+72 }=\frac{ x^{2}+8x-9 }{ x^{2}+17x+72 }\]

OpenStudy (anonymous):

you have to equate numerators since denominators are equal 2x^2+18x=x^2+8x-9 x^2+10x+9=0 (x+9)(x+1)=0 the equation has two roots: x=-9, x=-1

OpenStudy (anonymous):

Oh ok, I understand ! Thank you! :)

OpenStudy (anonymous):

@Hero I understand the problem & solution but my answer only requires one answer. I don't know what to put.

hero (hero):

You should try to understand the problem and the solution before inputting an answer.

OpenStudy (anonymous):

I do, but I get 2 answers when my online work only requires one for this question.

hero (hero):

When you get answers, you are supposed to "check" them to make sure they are true.

OpenStudy (anonymous):

My online work only gives me a certain amount of times to "check" my answer, then it'll lock the question whether it's wrong or right. This is my first time checking it, I only have 3 chances.

OpenStudy (anonymous):

@ganeshie8